OpenStudy (anonymous):

what is an equivalent representation of 5^-3 ?

OpenStudy (bookworm14):

do you need an expression or word form , etc?

OpenStudy (anonymous):

expression

OpenStudy (bookworm14):

well you can use \[\frac{ 1 }{ 125 } \] or 0.008 that i knw of- Hope this Helps- Bree :)

OpenStudy (bookworm14):

@kewlkatcutie

jhonyy9 (jhonyy9):

hope this will help you than you know a^(-2) = 1/(a^2) yes ?